Joe bought a pumpkin that cost 10 cents more per pound than his sister’s. Together, the two pumpkins weighed 20 pounds, but Joe’s pumpkin was heavier. Joe paid $3.52, and his sister paid $0.48. How many pounds did Joe’s pumpkin weigh?
Answer by richwmiller(17219)

price
j=10+s
weight
J=20-S
j*J=352
s*S=48
j = 22
s = 12
J = 16 joe's pumpkin weighs 16 pounds
S = 4